Benson, Inc., has sales of $45,530, costs of $14,550, depreciation expense of $3,350, and interest expense of $2,450. The tax rate is 25 percent.

What is the operating cash flow, or OCF? (Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1

Answer : $24,685

operating Cash flows = EBIT + Depreciation - Taxes

EBIT = Sales - Cost - Depraication = 45,530-14,550-3,350 = $27,630

TAX= (EBIT- Interest )* Tax Rate = (27,630-2,450)*25% = $6,295

Operating Cash flows = 27,630+3,350-6,295 = $24,685 (Answer)